# CHORDONOMICON: A Dataset of 666,000 Songs and their Chord Progressions

### v2

@@ Introduction
Numerous music datasets have been introduced incorporating a mix of various formats and metadata. They are utilized to address diverse tasks, including but not limited to genre classification (Dervakos, Kotsani, and Stamou 2021), music generation (Wang et al. 2020) and analysis (De Haas et al. 2011), chord recognition (Nadar, Abeßer, and Grollmisch 2019), key estimation (George, Mary, and George 2022), and more  [^2] https://www.music-ir.org/mirex/wiki/MIREX_HOME. In our work, we present a very large scale dataset featuring the symbolic representation of more than 666,000 contemporary music compositions through the use of music chords and chord progressions. We offer metadata for details such as genre, sub-genre, and release date. Additionally, we include structural information related to different parts of the music piece.
More specifically, in relation to Music Structural Analysis (MSA), there is a scarcity of datasets containing semantic labels (Wang, Hung, and Smith 2022) necessary for accurately defining distinct music components such as Verse and Chorus. It is notable that efforts in this field (Paulus and Klapuri 2009; Smith et al. 2011; Nieto et al. 2019) primarily emerged more than a decade ago, and the quantity of semantically labeled music tracks remains notably limited (around 900-1500). In our research, we provide 2,679,123 structural labels for 403,923 music tracks, adhering to the terminology proposed in these earlier studies.
@@
This section delves into insights related to data collection and processing. Specifically, we provide detailed descriptions of the methodology used for data collection and curation, while at the same time explaining the process of data selection and exclusion. Table 1 summarizes the dataset’s key statistics, offering a quick overview as we get ready to explore the upcoming discussions. On average, each progression consists of 76.375 chords, with a median of 70 chords and a standard deviation of 54.52.
@@
The next step was to format the raw HTML files into the full chord progression of each song, collapsing repeating identical chords into a single chord (’A G G A’ became ’A G A’), removing songs that had less than four chords in total and removing any songs that had odd chord encoding (e.g. Cyrillic chord characters) and used UTF-8 encoding to make sure all chords are same format. Finally, this remaining dataset was curated by music experts to retain the chord progressions where they were confident about the terminology of the chords (removing songs that contained bizarre or musically incoherent chord symbolisms), which gave us the ability to accurately convert them into the Harte syntax and the syntax proposed by the FHO. The selection and conversion were performed manually and were supervised by music experts: we compiled a document comprising all encountered chords, which was then reviewed by music experts who aligned them with the Harte syntax only when they were confident about the accuracy of the chord (e.g., “SOL5” to “G:(*3)”, “hmin” to “B:min”, etc.). Eventually, we conclude into the final dataset of 667,858 unique, chordified, musical tracks.
@@
Spotify Web API [^3] https://developer.spotify.com/documentation/web-api was utilized for metadata collection. Specifically, we gathered information about the release date of the tracks and the musical genres associated with their respective artists. To verify their correctness, we made certain that the names of both the music track and the artist in our data match the corresponding information in Spotify data.
@@
Additionally, we established a classification task specific to rock music using tracks whose artists are categorized within the rock genre. In this context, we retained the least prevalent rock genre associated with each track, focusing on genres with at least 100 occurrences. We’ve identified 83 distinct rock sub-genres. This task is considered a multi-class classification problem but is quite difficult due to the nuanced similarities between rock sub-genres.
Figure 1: Dataset genre frequency
Figure 2: Dataset part distribution
@@
No. of Tracks 667,858
No. of Chords 51,007,915
No. of Unique Chords 736
No. of Unique Inverted Chords 3,526
No. of Tracks with Parts 403,923
No. of Parts 2,679,123
No. of Tracks with Release Date 438,809
No. of Tracks with Spotify Genres 364,273
No. of Tracks with Main Genre 277,925
No. of Tracks with Rock Sub-Genre 121,411
@@
Our dataset comprises 51,007,915 chords, with 1,437,712 being inverted. Their types are shown in Table 2, while the number of their notes in Table 3. Figure 4 illustrates the chords’ distribution of our dataset. It’s evident that natural triads, especially the “G” chord, stand out as the most frequently occurring chords. This prevalence is likely due to the simplicity favored in modern music, which dominates our dataset, and the widespread use of capos. Additionally, since these chords are user-generated, they often reflect simplified interpretations of tracks, as non-professional transcribers may reduce complex chords to triads, even if the original music features more intricate harmonies.
@@
Major 36,331,633
Minor 13,104,882
Suspended 721,086
Power 717,252
Half-diminished 105,296
Augmented 27,400
Diminished 366
@@
Three 45,477,125
Four 4,457,563
Two 717,252
Five 285,140
Six 41,552
Seven 29,283
@@ Conclusion
1. User-Generated Content Ownership: UG’s TOS explicitly states that “Ultimate Guitar does not claim any ownership rights in User Generated Content that you transmit, submit, display or publish on, through or in connection with the Service”. And further clarifies that “User Generated Content includes, without limitation, tablatures (text or electronic)…”. This clause clearly delineates the ownership status of the content we’re collecting. 2. Non-Copyrightable Elements: Legal precedents consistently support the notion that chord progressions are not copyrightable. Cases such as Granite Music Corp. v. United Artists Corp. (1977) and Swirsky v. Carey (2004) have established that chord progressions are fundamental building blocks of music, akin to ”common musical property”, since many songs share the same or similar chord progressions. This legal stance corroborates that our dataset solely comprises of non-copyrightable elements. 3. Data Processing: To further ensure ethical and legal compliance, our data collection process excludes copyrightable elements such as lyrics, song titles, and artist names. We retain only the chord progressions and, where available, the associated song structure information (e.g., intro, verse, chorus), release year and genre of the song’s artist(s).

### v3

@@ Introduction
Numerous music datasets have been introduced incorporating a mix of various formats and metadata. They are utilized to address diverse tasks, including but not limited to genre classification (Dervakos, Kotsani, and Stamou 2021), music generation (Wang et al. 2020) and analysis (De Haas et al. 2011), chord recognition (Nadar, Abeßer, and Grollmisch 2019), key estimation (George, Mary, and George 2022), and more  [^2] https://www.music-ir.org/mirex/wiki/MIREX_HOME. In our work, we present a very large scale dataset featuring the symbolic representation of more than 666,000 contemporary music compositions through the use of music chords and chord progressions. We offer metadata for details such as genre, sub-genre, and release date. Additionally, we include structural information related to different parts of the music piece and Spotify IDs for tracks and artists.
More specifically, in relation to Music Structural Analysis (MSA), there is a scarcity of datasets containing semantic labels (Wang, Hung, and Smith 2022) necessary for accurately defining distinct music components such as Verse and Chorus. It is notable that efforts in this field (Paulus and Klapuri 2009; Smith et al. 2011; Nieto et al. 2019) primarily emerged more than a decade ago, and the quantity of semantically labeled music tracks remains notably limited (around 900-1500). In our research, we provide 2,670,457 structural labels for 397,580 music tracks, adhering to the terminology proposed in these earlier studies.
@@
This section delves into insights related to data collection and processing. Specifically, we provide detailed descriptions of the methodology used for data collection and curation, while at the same time explaining the process of data selection and exclusion. Table 1 summarizes the dataset’s key statistics, offering a quick overview as we get ready to explore the upcoming discussions. On average, each progression consists of 76.48 chords, with a median of 71 chords and a standard deviation of 54.59.
@@
The next step was to format the raw HTML files into the full chord progression of each song, collapsing repeating identical chords into a single chord (’A G G A’ became ’A G A’), removing songs that had less than four chords in total and removing any songs that had odd chord encoding (e.g. Cyrillic chord characters) and used UTF-8 encoding to make sure all chords are same format. Finally, this remaining dataset was curated by music experts to retain the chord progressions where they were confident about the terminology of the chords (removing songs that contained bizarre or musically incoherent chord symbolisms), which gave us the ability to accurately convert them into the Harte syntax and the syntax proposed by the FHO. The selection and conversion were performed manually and were supervised by music experts: we compiled a document comprising all encountered chords, which was then reviewed by music experts who aligned them with the Harte syntax only when they were confident about the accuracy of the chord (e.g., “SOL5” to “G:(*3)”, “hmin” to “B:min”, etc.). Eventually, we conclude into the final dataset of 679,807 unique, chordified, musical tracks.
@@
Spotify Web API [^3] https://developer.spotify.com/documentation/web-api was utilized for metadata collection. Specifically, we gathered information about the release date of the tracks and the musical genres associated with their respective artists. To verify their correctness, we made certain that the names of both the music track and the artist in our data match the corresponding information in Spotify data. We also provide the Spotify IDs of tracks and artists where available.
@@
Additionally, we established a classification task specific to rock music using tracks whose artists are categorized within the rock genre. In this context, we retained the least prevalent rock genre associated with each track, focusing on genres with at least 100 occurrences. We’ve identified 179 distinct rock sub-genres. This task is considered a multi-class classification problem but is quite difficult due to the nuanced similarities between rock sub-genres.
Figure 1: Dataset genres’ distribution
Figure 2: Dataset parts’ distribution
@@
No. of Tracks 679,807
No. of Chords 51,994,634
No. of Unique Chords 749
No. of Unique Inverted Chords 3,577
No. of Tracks with Parts 397,580
No. of Parts 2,670,457
No. of Tracks with Release Date 422,181
No. of Tracks with Spotify Genres 429,753
No. of Tracks with Main Genre 352,111
No. of Tracks with Rock Sub-Genre 145,218
No. of Tracks with Spotify Track ID 440,284
No. of Tracks with Spotify Artist ID 510,986
@@
Our dataset comprises 51,994,634 chords, with 1,464,392 being inverted. Their types are shown in Table 2, while the number of their notes in Table 3. Figure 4 illustrates the chords’ distribution of our dataset. It’s evident that natural triads, especially the “G” chord, stand out as the most frequently occurring chords. This prevalence is likely due to the simplicity favored in modern music, which dominates our dataset, and the widespread use of capos. Additionally, since these chords are user-generated, they often reflect simplified interpretations of tracks, as non-professional transcribers may reduce complex chords to triads, even if the original music features more intricate harmonies.
@@
Major 37,031,150
Minor 13,368,240
Suspended 731,520
Power 730,443
Half-diminished 105,281
Augmented 27,627
Diminished 373
@@
Three 46,529,077
Four 4,466,488
Two 730,443
Five 197,732
Six 41,738
Seven 29,156
@@ Conclusion
1. User-Generated Content Ownership: UG’s TOS explicitly states that “Ultimate Guitar does not claim any ownership rights in User Generated Content that you transmit, submit, display or publish on, through or in connection with the Service”. And further clarifies that “User Generated Content includes, without limitation, tablatures (text or electronic)…”. This clause clearly delineates the ownership status of the content we’re collecting. 2. Non-Copyrightable Elements: Legal precedents consistently support the notion that chord progressions are not copyrightable. Cases such as Granite Music Corp. v. United Artists Corp. (1977) and Swirsky v. Carey (2004) have established that chord progressions are fundamental building blocks of music, akin to ”common musical property”, since many songs share the same or similar chord progressions. This legal stance corroborates that our dataset solely comprises of non-copyrightable elements. 3. Data Processing: To further ensure ethical and legal compliance, our data collection process excludes copyrightable elements such as lyrics, song titles, and artist names. We retain only the chord progressions, the associated song structure information (e.g., intro, verse, chorus), release year, the genres of the relevant artist and Spotify IDs of track and artist (non-copyrightable alphanumeric strings).